OpenStudy (anonymous):

Find two geometric means between 7 and 875.

OpenStudy (anonymous):

35 and 175

OpenStudy (anonymous):

\[\frac{ x }{ 7 }=\frac{ y }{ x }=\frac{ 875 }{ y }\]

OpenStudy (anonymous):

\[y^2=875x\] and \[7y=x^2\] \[y=\frac{ x^2 }{ 7 }\] \[\frac{ x^4 }{ 49 }=875x\]

OpenStudy (anonymous):

\[x^3=42875=>x=35\]

OpenStudy (anonymous):

and u can easily see the common ratio is 5, as as the sequence we took was 7,x,y,875 so, x=35, so y=175!